OpenStudy (anonymous):

how many moles of h2o are produced when 64.0 g c2h2 burns in oxygen

OpenStudy (anonymous):

1st of all find the moles of C2H2 i.e moles = mass / m.mass = 64/26 = 2.45 2C2H2 + 5O2 ---> 2H20 + 4CO2 2 moles of C2H2 form = 2 moles of H2O 2.45 moles of C2H2 form = 2/2 X 2.45 = 2.45 moles of H2O